OpenStudy (alexh107):

Find the derivative of f(x) = 5/x at x = -1.

zepdrix (zepdrix):

Hey Alex, recall a helpful exponent rule,\[\large\rm \frac{1}{x}=x^{-1}\]

zepdrix (zepdrix):

Applying this rule will allow us to use our power rule for differentiation.

OpenStudy (alexh107):

So this would be x^-5?

zepdrix (zepdrix):

No but good guess.\[\large\rm \frac{5}{x}=5\cdot\frac{1}{x}\]

zepdrix (zepdrix):

The 5 can just simply come down in front. And then you can apply the rule to the x-portion sitting next to it.

OpenStudy (alexh107):

Is it -1?

zepdrix (zepdrix):

\[\large\rm f(x)=\frac5x\qquad\qquad\qquad\to\qquad\qquad\qquad f(x)=5x^{-1}\]Ok yes, good. That's how we bring the x up to the numerator, we apply a -1 power to it. Now it's ready for power rule. Remember how to do that?

OpenStudy (alexh107):

-1 * 5 = -5?

zepdrix (zepdrix):

I don't quite understand.